# Multi-FAct: Assessing Multilingual LLMs’ Multi-Regional Knowledge using FActScore

### v2

@@ Abstract
Large Language Models (LLMs) are prone to factuality hallucination, generating text that contradicts established knowledge. While extensive research has addressed this in English, little is known about multilingual LLMs. This paper systematically evaluates multilingual LLMs’ factual accuracy across languages and geographic regions. We introduce a novel pipeline for multilingual factuality evaluation, adapting FActScore (Min et al., 2023) for diverse languages. Our analysis across nine languages reveals that English consistently outperforms others in factual accuracy and quantity of generated facts. Furthermore, multilingual models demonstrate a bias towards factual information from Western continents. These findings highlight the need for improved multilingual factuality assessment and underscore geographical biases in LLMs’ fact generation [^1] Source code and data available at: https://github.com/sheikhshafayat/multi-fact.
@@ 1 Introduction
Large Language Models (LLMs) are susceptible to factuality hallucination, a phenomenon in which the generated text contradicts established world knowledge (Huang et al., 2023; Zhang et al., 2023). Despite extensive research focusing on hallucination and factuality of LLMs in free-form generation, these endeavors have predominantly concentrated on English (Huang et al., 2023; Min et al., 2023; Mishra et al., 2024). Consequently, there exists a notable gap in our comprehension of the factual accuracy of LLMs when producing content in various languages. As highlighted by Kang et al. (2024), current metrics for detecting hallucination are inadequate in multilingual settings. The extent to which LLMs exhibit factual hallucination remains unclear across different languages. Similar to other capabilities, there may be a discernible decline in performance when LLMs are tasked with non-English contexts (Ahuja et al., 2023; Bang et al., 2023).
(a) Factuality Heatmap of GPT-3.5 Biography Generation in English
(b) Factuality Heatmap of GPT-3.5 Biography Generation in Korean
Figure 1: While GPT-3.5 in English knows about the presidents of all countries almost uniformly, its knowledge in Korean is much smaller and mostly limited to East Asia and North America. Here color represents FActScore (bluer is better).
@@
In this paper, we address this gap by systematically evaluating the factual accuracy of multilingual LLMs across different languages and geographic regions. We explore the following research questions: R1: Do multilingual models, such as ChatGPT, exhibit uniform factual accuracy across all languages in free-form generation? R2: Do these models demonstrate higher factual precision for content that aligns closely with the language (e.g., discussing a president of a country whose official language is used for generation)?
To address these questions, we introduce a novel pipeline tailored for evaluating factuality in a multilingual setting. Our approach first adapts the FActScore (Min et al., 2023) to accommodate multiple languages, and we make this pipeline openly accessible as open-source. Using this pipeline, we then assess the factual accuracy of nine languages in biography generation tasks, ensuring geographical diversity through a curated list of subjects.
@@
Through our analysis, we observe two significant findings. First, English consistently maintains an advantage in both factual accuracy and the quantity of generated facts compared to other languages when generating identical content. Second, content produced by multilingual language models tends to exhibit a stronger performance for factual information originating from Western regions, such as America and Europe, across the languages. These results emphasize the need for enhanced assessment methods in evaluating multilingual factual accuracy and underscore the geographical biases in the fact generation capabilities of LLMs.
2 Methodology
2.1 Task Selection
Min et al. (2023) select biography generation as their task because biographies contain factual information readily verifiable from existing sources like Wikipedia, and confirm through human evaluation that English Wikipedia is suitable for biography verification tasks.
Topic Selection
We select “National Leaders of Each Countr” as the topic for factuality measurement, as it is fairly common across various cultures and languages to have well-known and well-described national leaders such as presidents and prime ministers. Specifically, we choose to focus on the president or head of state in the year 2015. This year is randomly selected from those prior to the emergence of LLMs. To determine our samples of the topic, we select countries by compiling a list of the 20 most populous nations from each continent. [^2] America, Asia, Europe, and Africa; America consisted of North and South America; Australia and New Zealand were also grouped with America. The lists of countries and their corresponding leaders used in our experiments are detailed in Table 2 of the Appendix. We choose a total of 9 languages, divided among high, middle and low resource languages: English, German, French, Spanish, Arabic, Swahili, Chinese, Korean and Bengali. We make sure to include at least three languages that are widely spoken in each continent. [^3] Africa: Arabic, Swahili, and French; Europe: Spanish, French, German; Asia: Korean, Bengali, Chinese; America: Spanish and French. English is widely spoken all over the world, and kept as the baseline.
Figure 2: Pipeline of our methodology. For details of step three, fact verification, please refer to the original FActScore Min et al. (2023).
Figure 2: Pipeline of our methodology. For details of step three, fact verification, please refer to the original FActScore Min et al. (2023).
2.2 Multi-FAct Pipeline
We introduce a novel pipeline for automatically measuring factuality of a multilingual LLM for multi-regional knowledge. This pipeline evaluates the biographies of presidents or heads of 80 countries across 9 languages.
@@ 2.2 Multi-FAct Pipeline
The pipeline as shown in Figure 2 is structured into three main stages: 1) Obtaining multilingual generations (§ 2.3), 2) Translating these facts into English using GPT-3.5 (§ 2.4), 3) Measuring Factuality(§ 2.5).
@@ 2.3 Generating Facts
Following  Min et al. (2023), we prompt the model to generate a biography of designated topics.
@@
We use GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-1106-preview) for the generation of facts. All experiments were carried out between January and February 2024. The model temperature was set at 1.0.
@@
We translate the following prompt into 9 languages: Write a biography of {name}. The prompts were translated by GPT-4 and verified by native speakers’ manual evaluation. We also transliterate each person’s name into corresponding languages and validated by native speakers.
@@ 2.4 Translating Generated Facts
To compare the biographies written in the eight non-English languages, we translate the generated content from the original languages to English using GPT-3.5, specifically, gpt-3.5-turbo-0125. In principle, any translation model that can generate coherent translation would work. We do not use commercially available machine translators like Google Translate because manual evaluation revealed that they can not maintain the consistent gender of the person throughout the whole text.
We translate the generation and verify facts in English, rather than doing fact verification in corresponding languages with respective Wikipedia articles, for several reasons. Firstly, Wikipedia differs in size and scope for each language Wikipedia contributors (2024), which makes it difficult to compare cross-lingual factuality if the knowledge base is different. Moreover, some important steps in the original FActScore pipeline work best in English (e.g., RAG, NPM are designed primarily for English). We, however, do some quantitative analysis on the effect of LLM-based translation on FActScore evaluation and find that translation by GPT-3.5 does not affect FActScore of texts significantly (see appendix A.2).
@@ 2.5 Measuring Factuality
To evaluate the accuracy of the generated model’s response M, we use the FActScore metric. This involves breaking down the translated generation into atomic facts—short sentences conveying a single piece of information, as defined in the FActScore paper. The accuracy of these atomic facts, denoted as A, is verified against the corresponding English Wikipedia article, serving as the reference knowledge source C.
@@
We replace all proprietary models of the original FActScore pipeline with open-source models, ensuring cost-effectiveness while maintaining the quality of our system. We decompose LLM responses M into atomic facts using Mistral-7B Jiang et al. (2023). The verification step also uses Mistral-7B along with RAG and NPM Min et al. (2022) for the best performance. The details statistics can be found in Appendix A.2.
3 Results
@@
Figure 3: FActScore for each language of GPT-3.5 and GPT-4 for each language.
Figure 4: Number of correct and hallucinated facts by GPT-3.5 and GPT-4 for each language. The number on top of each bar is the FActScore . Note that in some languages ChatGPT generates shorter responses, means even though they might be generating not very good responses (or refusing to respond), their FActScore is higher.
3.1 Does the Factuality Differ Across Languages?
@@
Figure 3 displays the outcomes of measuring factuality across various languages using FActScore as a metric, uncovering significant variations among languages.
English, Spanish, French, German, and Swahili exhibit notably higher FActScore for both GPT-3.5 and GPT-4. Figure 5 shows the distribution of FActScore by language. Note that even in GPT4, some languages have long tail distribution in FActScore .
(a) GPT-4
(b) GPT-3.5
Figure 5: FActScore Distributions by language between GPT-4 and GPT-3.5
@@
Figure 4 compares factuality across languages through the average number of correct and hallucinated facts, highlighting a significant gap in the quantity of facts generated between English and other languages. Remarkably, even with similar FActScore as shown in Figure 3, English surpasses other languages in generating a higher number of correct facts, thus delivering more useful responses.

### v3

# Multi-FAct: Assessing Factuality of Multilingual LLMs using FActScore

@@ Abstract
Evaluating the factuality of long-form large language model (LLM)-generated text is an important challenge. Recently there has been a surge of interest in factuality evaluation for English, but little is known about the factuality evaluation of multilingual LLMs, specially when it comes to long-form generation. We introduce a simple pipeline for multilingual factuality evaluation, by applying FActScore (Min et al., 2023) for diverse languages. In addition to evaluating multilingual factual generation, we evaluate the factual accuracy of long-form text generation in topics that reflect regional diversity. We also examine the feasibility of running the FActScore pipeline using non-English Wikipediaand provide comprehensive guidelines on multilingual factual evaluation for regionally diverse topics.
@@ 1 Introduction
Large Language Models (LLMs) are susceptible to factuality hallucination, a phenomenon in which the generated text contradicts established world knowledge (Huang et al., 2023; Zhang et al., 2023). Despite extensive research focusing on hallucination and factuality of LLMs in free-form generation, the previous works have predominantly studied English (Huang et al., 2023; Min et al., 2023; Mishra et al., 2024; Wei et al., 2024; Wang et al., 2023). Consequently, there exists a wide gap in our comprehension of the factual accuracy of LLMs when producing content in non-English languages. As highlighted by Kang et al. (2024), current metrics for detecting hallucination are inadequate in multilingual settings. The extent to which LLMs exhibit factual hallucination remains unclear across different languages. Similar to other capabilities, there may be a discernible decline in performance when LLMs are tasked with non-English contexts (Ahuja et al., 2023; Bang et al., 2023).
@@
In this paper, we address this gap by systematically evaluating the factual accuracy of multilingual LLMs across different languages and geographic regions. We explore the following research questions: R1: How can we effectively evaluate the factuality of multilingual LLMs in free-form text generation? R2: How do multilingual models’ free-form answers compare in factual accuracy across languages and geographically contextualized questions?
To address these questions, we introduce Multi-FAct , a simple pipeline tailored for evaluating factuality in a multilingual context. Leveraging open-source models, we adapt the FActScore (Min et al., 2023) for multiple languages and validate it using human annotations from Min et al. (2023). Furthermore, we explore the use of non-English Wikipedia content to augment factuality evaluation, finding that while English Wikipedia remains a reliable source, concatenating articles from multiple non-English languages can sometimes provide estimates close to English Wikipedia.
@@
We conduct an evaluation of the factual accuracy of recent proprietary multilingual LLMs in nine languages using biography generation task. We ensure geographical diversity by curating subjects from various regions. Our analysis reveals two significant findings. First, English consistently maintains an advantage in both factual accuracy and the quantity of generated facts compared to other languages. Second, content produced by multilingual language models exhibits better factual accuracy about North America and Europe across the languages.
Our contributions are as follows:
  • We propose a novel pipeline Multi-FAct , that applies FActScore in a multilingual setting, and demonstrates the feasibility of conducting factuality evaluation of long-form generation in the multilingual biography task.
  • We explore the potential and limitations of using non-English resources, particularly non-English Wikipedia articles, for factuality evaluation of long-form generation, showing that reasonably large non-English Wikipedia articles can also serve as a viable source.
  • We introduce Multi-FAct as a tool to investigate cultural and geographic biases in LLMs, offering insights into how these biases manifest in LLM generated content.
@@
Figure 1: Our Multi-FAct Pipieline. The pipeline is structured into three main stages: 1) Obtaining multilingual generations, 2) Translating these facts into English using GPT-3.5, and 3) Measuring factuality, by breaking them down into smaller atomic facts and then verifying them by asking LLMs with context provided.
2 Related Work
LLM Factuality Evaluation
The assessment of factual accuracy in natural language texts predates the advent of LLMs (Guo et al., 2022). Our Multi-FAct pipeline follows the previous fact evaluation research, which involves validating claims based on external resource such as Wikipedia article (Thorne et al., 2018; Krishna et al., 2022; Zhong et al., 2019) or Google search (Chern et al., 2023; Wei et al., 2024). Our work is inspired by and based on Min et al. (2023) which proposes FActScore to measure the factuality of a long-form generated output of LLMs by breaking it down into atomic facts. Another line of work focuses on evaluating factual accuracy solely based on model output or internal states, eliminating the need for external knowledge bases like Wikipedia (Manakul et al., 2023; Azaria & Mitchell, 2023; Dhuliawala et al., 2023). While such methods offer resource efficiency, they often sacrifice accuracy compared to approaches leveraging external resources for factuality evaluation.
Multilingual Factuality Evaluation
To evaluate factual accuracy of multilingual generation, Gupta & Srikumar (2021) presents the X-Fact dataset. Similarly, Thorne et al. (2018) presents frameworks for multilingual fact-extraction and verification evaluation. Additionally, Aharoni et al. (2022) introduces mFACE, a dataset aimed at evaluating multilingual factual consistency. While these papers advance the understanding of multilingual factuality, they focus on the assessment of factual accuracy in existing articles or benchmarks rather than examining the factual accuracy in the long-form generation of LLMs. Recently, Kang et al. (2024) explores the efficacy of existing metrics designed to detect hallucinations of LLM-generations in English in a multilingual setting, showing that the English-based metrics fall short in multilingual contexts. Our study extends the endeavor of developing a factuality evaluation metric that can be applied in multiple languages by introducing a Multi-Fact pipeline.
Geo-culture biases of LLMs
Research shows LLMs exhibit cultural and linguistic bias (Hovy & Yang, 2021; Cao et al., 2023; Hershcovich et al., 2022; Huang & Yang, 2023; Jin et al., 2023; Naous et al., 2023). A recent paper conducts a true-false evaluation task that shows GPT models are more accurate for facts about the Global North compared to the Global South (Mirza et al., 2024), consistent with our findings. We conduct a much more comprehensive study with long-form generation of geographically diverse information in multiple languages.
3 Multi-FAct Pipeline
We introduce Multi-FAct, a novel pipeline for automatically measuring factuality of a multilingual LLM in a multilingual setting. Multi-FAct evaluates the biographies of multi-regional topics in multiple languages. We choose the biography generation task because biographies are suitable for being decomposed into verifiable and independent "atomic" facts (Min et al. (2023)).
The Multi-FAct pipeline is structured into three main steps: 1) Obtaining multilingual generations (§ 3.1), 2) Translating these facts into English using GPT-3.5 (§ 3.2), and 3) Measuring factuality of the translated facts (§ 3.3).
@@ 3.1 Generating Facts
Most of the results in this paper use GPT4 (gpt-4-1106-preview) and GPT3.5 (gpt-3.5-turbo-0613) models because at the time of writing this paper, those were the only two strong multilingual models [^1] We also provide FActScore estimation for some newer models in Appendix C. for LLM multilingual biography generation. All experiments are conducted between January and March 2024222Except for GPT4o and GPT4o-mini in the Appendix., and the model temperatures are set to their defaults for respective models.
Language
We choose a total of nine languages: English, German, French, Spanish, Arabic, Swahili, Chinese, Korean, and Bengali. We carefully choose languages to represent the level of existing resources–high, medium, and low, and to represent diverse regions–Africa, Europe, Asia, and America  [^3] Africa: Arabic, Swahili, and French; Europe: Spanish, French, German; Asia: Korean, Bengali, Chinese; America: Spanish and French. English is widely spoken all over the world, and kept as the baseline..
@@
We use the following prompt: Write a biography of {name}. We translate the prompt with GPT-4 into the eight non-English langauges, then we ask the native speakers to verify the translation. We also transliterate each person’s name into corresponding languages either using Wikipedia crosslinking or GPT-4, of which we take a subset and the native speakers verify that almost all the transliterations are correct.
@@ 3.2 Translating Generated Facts
To compare the biographies written in the eight non-English languages, we translate the generated content from the original languages to English using gpt-3.5-turbo-0125. We do not use commercially available machine translation systems, such as Google Translate, based on the preliminary result that they do not maintain the consistent gender of the person throughout the text [^4] However, we do provide an analysis of what happens when Google Translation is used instead of GPT3.5 in Appendix C..
We translate the generation and verify facts in English, rather than doing fact verification in corresponding languages with respective Wikipedia articles for multiple reasons. First, Wikipedia differs in size and scope for each language (Wikipedia contributors, 2024), which makes it difficult to compare cross-lingual factuality if the knowledge base is different. Additionally, key components of the original FActScore pipeline, such as RAG and NPM, are optimized for English and not available in other languages. Our quantitative analysis of LLM-based translation’s impact on FActScore evaluation reveals that GPT-3.5’s translation minimally influences the FActScore of texts (See § 3.4).
@@ 3.3 Measuring Factuality
To evaluate the accuracy of the generated model’s response M, we use the FActScore metric. This involves breaking down the translated generation into atomic facts—short sentences, each conveying a single piece of information, as defined in Min et al. (2023). We denote the set of atomic facts as A, and we measure the accuracy of these atomic facts with the corresponding English Wikipedia article serving as the reference knowledge source C.
@@
We replace all proprietary models of the original FActScore pipeline with open-source models, ensuring cost-effectiveness while maintaining the quality of our system. We decompose LLM responses M into atomic facts using Mistral-7B Jiang et al. (2023). The verification step also uses Mistral-7B along with RAG and NPM for the best performance. We set the NPM threshold at 0.3.
3.4 Reliability of Multi-FAct
3.4.1 Replication of original FActScore
We conduct a comparative analysis using our implementation, which includes a Mistral 7B model, instead of InstructGPT for atomic break-down, against the outcomes presented in the Min et al. (2023), using a subset of human-annotated factual data generated by ChatGPT (See Table 1). A comparison with other baselines and different parts of the Multi-FAct pipeline for different languages is provided in appendix A.
Metric Human FS FS (ours) Error
Value 0.626 $\pm$ 0.238 0.640 $\pm$ 0.127 -0.014 $\pm$ 0.128
Table 1: Summary of Replication. Error refers to the difference between human FActScore and FActScore determined by our method. Standard deviation is reported alongside each metric. Note that the Human FActScore value is slightly different from what is provided in Min et al. (2023) as this value is averaged over our subset (44 samples).
3.4.2 Effect of GPT-3.5 Translation in Multi-FAct
As our pipeline crucially relies on the GPT-3.5 translation (Step 3 in figure 1), it is imperative to check whether the translation step affects the automated estimation of FActScore in the multilingual setting. We evaluate the effect of GPT-3.5 translation based on the premise that the factual content of a text should remain consistent across translations into different languages. In other words, if the factuality of an English text is established, its translated version in another language should exhibit the same level of factuality.
Language Correlation Mean error Standard Deviation
English 0.84 -0.014 0.13
German 0.83 -0.007 0.13
French 0.80 -0.004 0.14
Spanish 0.81 -0.014 0.14
Swahili 0.81 -0.015 0.13
Arabic 0.82 0.021 0.14
Chinese 0.81 0.040 0.14
Korean 0.80 0.009 0.14
Bengali 0.84 0.027 0.13
Table 2: FActScore evaluation across languages. English refers to FActScore evaluation on original human annotation data provided in English; other languages refer to Multi-FAct evaluation on GPT-4 translation of the human annotation. Correlation refers to the correlation of individual FActScore with human-annotated FActScore provided in Min et al. (2023).
To simulate and test the reliability of Multi-FAct with respect to translation, we first take the human-annotated examples provided in the original work (Min et al., 2023) and translate them into the eight non-English languages using GPT-4555We use GPT-4 as a previous study (Jiao et al., 2023) demonstrated GPT-4 to be a good translator.. Then, we apply our multilingual FActScore pipeline (Figure 1) on the translated texts [^6] We did not use GPT-4 for back-translation in Multi-FAct pipeline to reduce API costs.. Table 2 indicates that the execution of Multi-FAct on non-English data does not significantly degrade the estimation of FActScore, and the correlation between FActScore in different languages and human evaluation remains comparable. This outcome implies that adding a translation step to our pipeline does not significantly influence the estimation of FActScore .
4 Reference Sources in Multilingual Factuality Evaluation
(a) Number of total Wikipedia articles available per language (in millions).
(b) Number of active Wikipedia users per language (note the log scale).
(c) Number of original FActScore annotated topics that have Wikipedia articles in corresponding languages.
Figure 2: The Wikipedia size distribution for languages in this study. (c) shows the number of human-annotated examples that also have representation in the corresponding Wikipedia. Note that Wikipedia size differs widely across languages, which usually means non-English Wikipedia articles are not as comprehensive as English Wikipedia articles for fact verification.
We evaluate the factual accuracy of LLM-generated biographies against English Wikipedia for its reliability and comprehensive coverage of information about individuals, following prior works (Petroni et al., 2020; Chen et al., 2017; Min et al., 2023). However, we need to analyze the reliability of non-English Wikipedia articles as references. While the Multi-FAct pipeline is effective for topics with established English Wikipedia articles, its applicability may be limited because many regional topics (e.g., local politicians) lack English Wikipedia articles. If we cannot extend our Multi-FAct pipeline to these topics, it would be a major obstacle to ensuring LLM factuality with geographical diversity. However, as figure 2 illustrates, languages included in our study exhibit a wide range of variability in their representation within Wikipedia.
In this section, we explore the following question: How should we measure multilingual factuality, when we cannot find a sufficient golden source of reference in English? First, we examine whether simple translation of Wikipedia articles from languages other than English would sustain the quality (§ 4.1). Then, we reveal that the "length of articles" is one factor contributing to performance degradation when non-English Wikipedia is used as fact verification ground truth corpus, and we propose a new method of concatenating articles from different languages to enhance the comprehensiveness of the reference (§ 4.2).
4.1 Using Translated Wikipedia Articles as Reference
We first test whether we can simply translate the Wikipedia articles written in eight non-English languages into English then replace the English Wikipedia articles within Multi-FAct pipeline with the translations. Using GPT-3.5, we translate the respective Wikipedia articles into English and conduct factual evaluation in English. We chose 157 individuals as topics for whom human-annotated data exists in Min et al. (2023), which served as the ground truth. We then compare this ground truth with the experimental results.
Figure 3(a) illustrates the errors for each language’s Wikipedia against ground truth, and we include English Wikipedia corpus as the baseline. The results show a near-zero error rate for English, and lower error rates in French, Spanish, and German. We observe relatively higher error rates for Bengali, Korean, Chinese, and Arabic. Note that the size and coverage of Wikipedia varies across languages, thus the number of samples used in the experiments also varies by language, as indicated by $n$ in the figures.
4.2 Effect of Reference Length
(a) FActScore estimation error when using different Wikipedia corpus as ground truth.
(b) FActScore estimation error scales linearly with the length of the Wikipedia article.
Figure 3: Effect of using non-English Wikipedia as knowledge corpus for FActScore estimation. The x-axis in subfigure 3(b) represents the word count (in English) after translation. As the length of the non-English Wikipedia article increases, its utility for evaluating factuality increases.
(a) Multi-FAct estimation for persons with ‘large’ Wikipedia articles in corresponding language.
(b) Ensembling multiple non-English Wikipedias work better than single language Wikipedia.
Figure 4: Effect of corpus size on FActScore Estimation. Figure 4(a) shows that comprehensive non-English sources might be used for factuality evaluation after translating into English.
We hypothesize that non-English Wikipedias result in higher errors compared to English Wikipedia due to the lower coverage of topics, as evidenced by the shorter length of non-English articles. Figure 3(b) shows that there is a linear relationship between FActScore estimation error and the word count of Wikipedia articles, indicating that longer Wikipedia articles tend to serve better as references for factuality evaluation.
To further test the linear relationship observed in Figure 3(b), we select 100 individuals from each of our studied languages who have both English Wikipedia articles and substantially long Wikipedia articles in their regional language. [^7] The detailed sampling method is provided in appendix C.2 We generate biographies for these 100 individuals in English and evaluate these biographies against both the English Wikipedia and the Wikipedia in the corresponding regional language. Since FActScore estimation for English is nearly perfect, we use it as a benchmark to estimate the error when using non-English articles as references. Figure 4(a) illustrates that the linear relationship between the length of Wikipedia articles and FActScore holds for longer articles as well.
Ensembling Wikipedia
Building upon the insights about the reference length discussed in § 4.1, this section investigates whether combining multiple non-English Wikipedias, referred to as “ensembling Wikipedias,” can enhance the accuracy of Multi-FAct measurements. In our experiments, we explore three cases: the union of wikis from Asian languages (Asian), the union of wikis from European languages (Euro), and the union of all wikis except English (All-Eng). As depicted in the Fig 3(a), among the two wikis with errors close to zero—namely, “Euro” and “All-Eng” cases—it is speculated that this could be attributed to their comprehensiveness or the high content overlap with the English Wikipedia. In the case of “Asian,” an error rate of 0.2 is observed, representing a decrease compared to using only one wiki of Asian language, and even approaching the performance observed for European languages (Spanish, French, and German).
Our findings suggest that even for the same topic, the coverage of Wikipedia articles may vary across languages, thereby demonstrating that an “ensemble Wikipedia” comprising multiple articles can mitigate the performance gap between non-English Wikipedias and the English Wikipedia. However, the challenge of Wikipedia containing only a single, brief article remains unaddressed.
5 Factual Accuracy of Multilingual models
We use Multi-FAct to investigate the factual accuracy of multilingual LLMs in biography generation. We select “National Leaders of Each Country” as the topic for factuality measurement, as it is fairly common across various cultures and languages to have well-known and well-described national leaders such as presidents and prime ministers. Specifically, we choose to focus on the president or head of state in the year 2015. This year is randomly selected from those prior to the emergence of LLMs and also to make sure that at least some amount of data about the leader is online, which might not be true for very recent leaders. The lists of countries and their corresponding leaders used in our experiments are detailed in Table 8 of Appendix.
5.1 Does factuality differ across languages?
Figure 5: FActScore of GPT-3.5 and GPT-4 for each language. Note that GPT-4 makes a strong jump from GPT-3.5, especially in low-resource languages.
@@ 5.1 Does factuality differ across languages?
Figure 6: Number of correct and hallucinated facts by GPT-3.5 and GPT-4 for each language. The number on top of each bar is the FActScore .
@@
Figure 6 illustrates the factuality across various languages using FActScore as a metric, uncovering significant variations among languages. English, Spanish, French, German, and Swahili exhibit notably higher FActScore for both GPT-3.5 and GPT-4.
@@
Figure 6 compares factuality across languages through the average number of correct and hallucinated facts, highlighting a significant gap in the quantity of facts generated between English and other languages. Notably, even with similar FActScore as shown in Figure 6, English surpasses other languages in generating a higher number of correct facts, thus delivering more useful responses.
@@
5.2 Does the Factuality Geographically Differ Across Languages?
In this section, we explore whether the factuality of models varies with the geographic distribution of the language. Since GPT-4 performs better in low resource language (Figure 6), from here on, we conduct our analysis using only the outputs from GPT-4.
Next, we analyze the geographic distribution of the most factually accurate information across different languages. For each language, we identify the continents associated with top-20 FActScore . Our findings reveal a strong bias towards American and European topics across all languages studied. This pattern persists even in languages like geographicallly distant from these continents, such as Korean and Chinese. For Arabic, while the highest FActScore is linked to Africa, likely reflecting its geographical roots, American topics still rank second place. This trend highlights a Western-centric bias in the model’s factual content distribution across languages. The result figure and detailed analysis of geographical biases across more fine-grained subregions are presented in Appendix D.